OpenStudy (anonymous):

Simplify 11 cubed root of 6 end root plus 8 fourth root of 6 end root plus 2 cubed root of 6 end root minus 3 fourth root of 6

OpenStudy (jdoe0001):

\(\bf \large {11\sqrt[3]{6}+8\sqrt[4]{6}+2\sqrt[3]{6}-3\sqrt[4]{6}\qquad say\qquad \sqrt[3]{6}=a\qquad \sqrt[4]{6}=b\\ \quad \\ 11\sqrt[3]{6}+8\sqrt[4]{6}+2\sqrt[3]{6}-3\sqrt[4]{6}\implies 11a+8b+2a-3b}\) sum like terms

OpenStudy (anonymous):

ugh this isnt gonna work im so confused im just gonna get a touter